# On the Weight Spectrum of Rate-Compatible Polar Codes

### v2

@@ IV Number of minimum-weight codewords in QUP and Wang-Liu shortened polar codes
In this section, we calculate the number of polynomials with minimum weight in the decreasing monomial code with QUP pattern and Wang-Liu shortening pattern.
Due to the variability in the weight spectrum of different polynomials after puncturing, to facilitate a clear analysis, define $P_{f}(d,i)$ to be the number of codewords in $T_{P}(f,X_{i})$ with weight $d$. Firstly, we analyse the weight spectrum of $T_{P}(f,X_{i})$.
@@ IV-A Weight spectrum of $T_{P}(f,X_{i})$
For polynomial $f=x_{i_{1}}\cdots x_{i_{t}}$, where the variables are ordered such that $i_{1}<\dots<i_{t}$. We have $p\in T(f)$ satisfies $p_{[1,2^{i_{t}}]}=p_{[k2^{i_{t}+1}+1,(k+1)2^{i_{t}}]}$ for all $10\leq k\leq 2^{m-i_{t}}-$ 1. Moreover, $\text{wt}(p_{[s,s+2^{i_{t}}-1]})=\text{wt}(p)/2^{m-i_{t}}$ for all $11\leq s\leq 2^{m}-2^{i_{t}}+$ 1.
@@
Note that for any $p\in T(f)$, $x_{j}$ with $j>i_{t}$ does not appear in $p$. For any $1\leq l\leq 2^{i_{t}}$ and $10\leq k\leq 2^{m-i_{t}}-$ 1, denote $l=D(a_{1},\dots,a_{m})$ and $l+k2^{i_{t}}=D(b_{1},\dots,b_{m})$, then $a_{j}=b_{j}$ for $1\leq j\leq i_{t}$, so $p_{l}=p(a_{1},\dots,a_{m})=p(b_{1},\dots,b_{m})=p_{l+k2^{i_{t}}}$. Therefore, $p_{[1,2^{i_{t}}]}=p_{[k2^{i_{t}}+1,(k+1)2^{i_{t}}]}$, and it follows that $\text{wt}(p_{[s,s+2^{i_{t}}-1]})=\text{wt}(p)/2^{m-i_{t}}$ for all $11\leq s\leq 2^{m}-2^{i_{t}}+$ 1. ∎
@@
For polynomial $f=x_{i_{1}}\cdots x_{i_{t}}$, where the variables are ordered such that $i_{1}<\dots<i_{t}$. The affine automorphism, denoted as $\varphi$, which maps $x_{j}$ to $x_{j}+1$ for all $j<i_{t}$, while leaving $x_{j}$ for $j\geq i_{t}$ unchanged satisfies that $\text{ev}(\varphi(p))_{k}=\text{ev}(p)_{2^{i_{t}}+1-k}$ for all $p\in T(f)$ and $1\leq k\leq 2^{i_{t}}$.
@@
For a monomial $f=x_{i_{1}}\cdots x_{i_{t}}$ with $i_{1}<\dots<i_{t},t\geq 2$, denote $N_{f}(w,a)=|\{p\in T(f)\mid\text{wt}(p_{[1,a]})=w\}|$ to be the number of codewords with weight-$w$ at the first $a$ bits, clearly $P_{f}(w,a)=N_{f}(2^{m-t}-w,a)$. We have
@@
[…]
Here, $\alpha_{f}(s)=|\{1\leq j\leq t-1\mid i_{j}>s\}|$ is the number of $i_{1},\dots,i_{t-1}$ larger than $s$, $f^{(0)}=x_{i_{1}}\cdots x_{i_{t-1}}$ and $f^{(s)}=x_{i_{1}}\cdots x_{i_{t-1}}x_{s}$, where $s\in B(f,t)$.
@@
1) $a>2^{i_{t}}$. From Lemma 3, wt$(p_{[a-2^{i_{t}}+1,a]})=2^{i_{t}-t}$. Therefore $\text{wt}(p_{[1,a]})=w$ if and only if $\text{wt}(p_{[1,a-2^{i_{t}}]})=w-2^{i_{t}-t}$, i.e.,
@@
2) $2^{i_{t}-1}<a\leq 2^{i_{t}}$. Form Lemma 4, there exists some automorphism $\varphi$, such that $p^{\prime}=\varphi(p)$ satisfying $p^{\prime}_{k}=p_{2^{i_{t}}+1-k}$ for $1\leq k\leq 2^{i_{t}}$, so $wt(p_{[1,a]})=wt(p^{\prime}_{[2^{i_{t}}+1-a,2^{i_{t}}]})=2^{i_{t}-t}-wt(p^{\prime}_{[1,2^{i_{t}}-a]})$. Since $\varphi$ is bijective, the number of codewords in $T(f)$ with $w$ ones in the first $a$ bits is equal to the number of codewords in $T(f)$ with $2^{i_{t}-t}-w$ ones in the first $2^{i_{t}}-a$ bits, that is,
@@
Then $p_{[1,a]}=\rho(p)_{[1,a]}$ since $x_{i_{t}}=1$ at the first $2^{i_{t}-1}$ bits. Next, we will demonstrate that $\rho(p)\in T(g)$ for a suitable $g\preccurlyeq f$, the precise choice of $g$ depends on the value of $a_{i_{t},k},k\in B(f,t)$. This allows us to compute $N_{f}(w,a)$ in an iterative fashion, consider the following three cases:
@@
III) If $a_{i_{t},k}=1$ for some $k\in B(f,t)$, we divide the remaining polynomials in $T(f)$ according to their largest integer $s$ such that $a_{i_{t},s}=1$. For $s_{1}\in B(f,t)$ and $p_{1}\in T(f^{(s_{1})})$, define $H_{p_{1}}\triangleq\{p\in T(f)\mid p_{[1,2^{i_{t}-1}]}=(p_{1})_{[1,2^{i_{t}-1}]}\}$. Note that for any $p_{1}\in T(f^{(s_{1})})$ and $p_{2}\in T(f^{(s_{2})})$, $(p_{1})_{[1,2^{i_{t}-1}]}\neq(p_{2})_{[1,2^{i_{t}-1}]}$, otherwise, from Lemma 3, $p_{1}=p_{2}$ since $s_{1},s_{2}\leq i_{t}-1$, which is a contradiction. Therefore,
$H_{p_{1}}\cap H_{p_{2}}=\varnothing.$ (6)
Next, we are going to prove there are $|H_{p_{1}}|=2^{\alpha_{f}(s_{1})}$ for all $s_{1}\in B(f,t)$ and $p_{1}\in T(f^{(s_{1})})$.
Denote $h=x_{s_{1}}+\sum_{k\in B(f,t),k<s}a_{i_{t},k}x_{k}+a_{i_{t},0}+1$, then $p_{1}=h\prod_{j=1}^{t-1}(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0})\in T(f^{(s_{1})})$. Now for $i_{j}>s_{1}$, then $h(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0})=h(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0}+h+1)$. Thus, for all $\beta_{j}\in\{0,1\}$, $(x_{i_{t}}+h+1)\prod_{i_{j}<s}(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0})\prod_{i_{j}>s,j\leq t-1}(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0}+\beta_{j}(h+1))$ are equal to $p_{1}$ at the first $2^{i_{t}-1}$ bits, so
$|H_{p_{1}}|\geq 2^{\alpha_{f}(s_{1})}.$ (7)
We have already shown there are at least $2^{\alpha_{f}(s_{1})}$ polynomials in $T(f)$ equal to $p_{1}$ at the first $2^{i_{t}-1}$ bits. Next we claim that $|H_{p_{1}}|$ is exactly $2^{\alpha_{f}(s_{1})}$. For $s_{1},s_{2}\in B(f,t)$, assume $s_{2}<s_{1}$ and there does not exist $s_{2}<s_{3}<s_{1}$ such that $s_{3}\in B(f,t)$. Then $s_{3}\in F_{f}$, so $\alpha_{f}(s_{1})+s_{1}=\alpha_{f}(s_{2})+s_{2}+1$, and. Furthermore, if $s_{0}$ is the minimum integer in $B(f,t)$, then $\alpha_{f}(s_{0})+s_{0}=t$. Hence, if $s$ is the $k$-$th$ smallest integer in $B(f,t)$, $\alpha_{f}(s)+s=t+k-1$. Since $|B(f,t)|=i_{t}-t$, we have
@@
$\lambda_{f}-2\lambda_{f^{(0)}}$ $\geq\sum_{s\in B(f,t)}\sum_{p_{1}\in T(f^{(s)})}|H_{p_{1}}|$
@@
Here, the first inequality is from (6). Note that the numbers of polynomials in case I) and II) are both $\lambda_{f^{(0)}}$, which needs to be subtracted. The second inequality is from (7). The equality is from multiplying both side of (8) by $\lambda_{f^{(0)}}$. Therefore, $|H_{p_{1}}|=2^{\alpha_{f}(s_{1})}$ for all $s_{1}\in B(f,t)$ and $p_{1}\in T(f^{(s_{1})})$.
@@
In Table 1, we have $P_{f}(w,a)=N_{f}(8-w,a)=N_{f}(6-w,a-8)=P_{f}(w+2,a-8)$ for $8\leq a\leq 16$, which is from (5d). Also, we note that $P_{f}(w,a)=P_{f}(14-w,8-a)$ when $4\leq a\leq 8$ according to Equation (5c).
For $1\leq a\leq 4$ and $w\neq 8$, $P_{f}(w,a)$ is calculated by Equation (5). In fact, we know that the minimum codeword in $T(x_{2}x_{3})$ can be represented as
@@
When $1\leq a\leq 4$, $x_{3}=1$, so $f_{[1,4]}=g_{[1,4]}$, where
$p_{1}=(a_{3,1}x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}x_{1}+a_{2,0}).$
Now there are three cases for $p_{1}$:
Case I: $a_{3,1}=0,a_{3,0}=1$, then $p_{1}=0$.
Case II: $a_{3,1}=a_{3,0}=0$, then $p_{1}=x_{2}+a_{2,1}x_{1}+a_{2,0}\in T(x_{2})$.
Case III: $a_{3,1}=1$, then $p_{1}=(x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}x_{1}+a_{2,0})\in T(x_{2}x_{1})$. Note that $(x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}x_{1}+a_{2,0})=(x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}(a_{3,0}+1)+a_{2,0})$. One polynomial $p_{1}$ in $T(x_{2}x_{1})$ corresponds to two polynomials in $T(x_{1})$.
In conclusion, $N_{f}(w,a)=N_{x_{2}}(w,a)+2N_{x_{1}x_{2}}(w,a)$ for $1\leq a\leq 4$, and $P_{f}(w,a)=N_{f}(8-w,a)$. Finally, $P_{f}(8,a)=N_{f}(0,a)=16-\sum_{w=1}^{8}N_{f}(w,a)$ from Equation (5).
@@ IV-B Number of minimum-weight codewords in QUP polar codes
Building upon Lemma 5, we derive the weight spectrum of $T_{P}(f,X_{i})$ after puncturing. Next, we prove that a minimum-weight codeword after puncturing must also be a minimum-weight codeword under the mother code length, i.e., belong to $T_{P}(f,X_{i})$.
Let ${\bm{f}}_{N}^{(j)}$ be the $j$-$th$ row vector of $\bm{F}_{N}$. Note that the minimum weight of $\mathcal{C}(m,(\bm{0}_{1}^{j-1},1))$ is wt$({\bm{f}}_{N}^{(j)})$ [22, Corollary 1]. Also note that $\text{wt}({\bm{f}}_{N/2}^{(j)})=\text{wt}({\bm{f}}_{N}^{(j)})=\text{wt}({\bm{f}}_{N}^{(j+N/2)})/2$ for $21\leq j\leq N/$ 2.
@@
$({\bm{f}}_{N}^{(j)})_{[N]/X_{i}}$ is a minimum-weight codeword in $\mathcal{C}(m,(\bm{0}_{1}^{j-1},1),X_{i})$. Furthermore, the minimum-weight codewords in $\mathcal{C}(m,(\bm{0}_{1}^{j-1},1),X_{i})$ are also minimum-weight codewords under the mother code length.
@@
We prove the lemma by induction on $m$. The case $m=1$ can be proved directly. For the inductive step $m-1\to m$, we consider two cases:
1) $21\leq j\leq N/$ 2, i.e., ${\bm{f}}_{N}^{(j)}$ is in the top half of $\bm{F}_{N}$. Then $\mathcal{C}(m,(\bm{0}_{1}^{j-1},1),X_{i})=\{(\bm{u}_{[N/2]/X_{i}}\oplus\bm{v}_{[N/2]/X_{i}},\bm{v})\mid\bm{u}\in\mathcal{C}(m-1,(\bm{0}_{1}^{j-1},1),\bm{v}\in\mathbb{F}_{2}^{N/2}\}$. Then by induction hypothesis,
@@
Furthermore, if $(\bm{u}_{[N/2]/X_{i}}+\bm{v}_{[N/2]/X_{i}},\bm{v})$ is the minimum-weight codeword, then the equality in (IV-B) must hold. By inductive hypothesis, $\text{wt}(\bm{u}_{[N/2]/X_{i}})=\text{wt}(({\bm{f}}_{N/2}^{(j)})_{[N/2]/X_{i}})$ implies that $\bm{u}$ is the minimum-weight codeword under the mother code length $N/2$. Besides, $\text{wt}((\bm{u}_{[N/2]/X_{i}}\oplus\bm{v}_{[N/2]/X_{i}},\bm{v}))=\text{wt}(\bm{u}_{[N/2]/X_{i}})$ implies that $\text{wt}(\bm{v}_{X_{i}})=0$ and $\bm{v}_{[N/2]/X_{i}}$ is covered by $\bm{u}_{[N/2]/X_{i}}$, i.e, $v_{k}=1$ implies $u_{k}=1$ for any $k\in[N/2]/X_{i}$. Hence, $\bm{v}$ is covered by $\bm{u}$, $\text{wt}((\bm{u}+\bm{v},\bm{v}))=\text{wt}(\bm{u})=\text{wt}({\bm{f}}_{N/2}^{(j)})=\text{wt}({\bm{f}}_{N}^{(j)})$ and $(\bm{u}\oplus\bm{v},\bm{v})$ is the minimum-weight codeword under the mother code length $N$.
@@
Furthermore, if $(\bm{v}_{[N/2]/X_{i}},\bm{v})$ is the minimum-weight codeword, then the equality in (IV-B) must hold, that is, $\text{wt}(\bm{v})=\text{wt}({\bm{f}}_{N/2}^{(j)})$. Then $(\bm{v},\bm{v})$ is the minimum-weight codeword under the mother code length $N$.
In conclusion, the minimum-weight codewords in $C(\mathcal{I},X_{i})$ are also minimum-weight codewords under the mother code length, i.e., belong to some $T_{P}(f,X_{i})$. ∎
The number of minimum-weight codewords in QUP polar code then follows from Lemma 5 and Lemma 6.
@@
Suppose $C(\mathcal{I},X_{i})$ is a punctured decreasing $r$-th order polar code with length $N=2^{m}$ and puncturing pattern $X_{i}$, where $X_{i}$ is the first $i$ bits of the sequence $\bm{q}$. Denote $d$ to be the minimum positive integer such that $\sum_{f\in\mathcal{I}}P_{f}(d,X_{i})>0$. Then $d$ is the minimum weight of $C(\mathcal{I},X_{i})$ and $A_{d}(C_{P}(\mathcal{I},i))=\sum_{f\in\mathcal{I}}P_{f}(2^{m-r}-d,X_{i})$ and $A_{w}(C_{P}(\mathcal{I},i))\geq\sum_{f\in\mathcal{I}}P_{f}(2^{m-r}-w,X_{i})$ for $w>d$.
The Algorithm 2 concludes the procedure for calculating the number of minimum codewords in QUP polar code. In fact, the lower bound $\sum_{f\in\mathcal{I}}P_{f}(2^{m-r}-w,i)$ in Theorem 3 for $w$ larger than the minimum weight can also be calculated by Lemma 5.
Algorithm 2 Calculate the number of minimum codewords in QUP polar code
@@
As $m$ approaches infinity, the time complexity of Algorithm 2 is $O(N^{3})$.
@@
Since LTA transformation is the automorphism of decreasing polar codes, it will result in the same weight spectrum if the puncturing pattern is transformed by an LTA transformation. Consequently, we could know weight spectrum for a series of punctured decreasing polar codes.

### v3

@@ IV Number of minimum-weight codewords in QUP and Wang-Liu shortened polar codes
In this section, we calculate the number of polynomials with minimum weight in the decreasing monomial code with QUP and Wang-Liu shortening patterns.
Due to the variability in the weight spectrum of different polynomials after puncturing, we define $P_{f}(d,i)$ to be the number of codewords in $T_{P}(f,X_{i})$ with weight $d$ to facilitate a clear analysis. First, we analyse the weight spectrum of $T_{P}(f,X_{i})$.
@@ IV-A Weight spectrum of $T_{P}(f,X_{i})$
For polynomial $f=x_{i_{1}}\cdots x_{i_{t}}$ with $i_{1}<\dots<i_{t}$, $p\in T(f)$ satisfies $p_{[1,2^{i_{t}}]}=p_{[k2^{i_{t}+1}+1,(k+1)2^{i_{t}}]}$ for all $0\leq k\leq 2^{m-i_{t}}-1$. Moreover, $\text{wt}(p_{[s,s+2^{i_{t}}-1]})=\text{wt}(p)/2^{m-i_{t}}$ for all $1\leq s\leq 2^{m}-2^{i_{t}}+1$.
@@
Note that for any $p\in T(f)$, the variable $x_{j}$ with $j>i_{t}$ does not appear in $p$. For any $1\leq l\leq 2^{i_{t}}$ and $0\leq k\leq 2^{m-i_{t}}-1$, we denote $l=D(a_{1},\dots,a_{m})$ and $l+k2^{i_{t}}=D(b_{1},\dots,b_{m})$. Because $a_{j}=b_{j}$ for $1\leq j\leq i_{t}$, it follows that $p_{l}=p(a_{1},\dots,a_{m})=p(b_{1},\dots,b_{m})=p_{l+k2^{i_{t}}}$. Therefore, $p_{[1,2^{i_{t}}]}=p_{[k2^{i_{t}}+1,(k+1)2^{i_{t}}]}$, and $\text{wt}(p_{[s,s+2^{i_{t}}-1]})=\text{wt}(p)/2^{m-i_{t}}$ holds for all $1\leq s\leq 2^{m}-2^{i_{t}}+1$. ∎
@@
Let $f=x_{i_{1}}\cdots x_{i_{t}}$ with $i_{1}<\dots<i_{t}$. We define an affine automorphism $\varphi$ that maps $x_{j}$ to $x_{j}+1$ for all $j<i_{t}$ while leaving $x_{j}$ for $j\geq i_{t}$ unchanged. Then $\varphi$ satisfies that $\varphi(p)_{k}=p_{2^{i_{t}}+1-k}$ for all $p\in T(f)$ and $1\leq k\leq 2^{i_{t}}$.
@@
For a monomial $f=x_{i_{1}}\cdots x_{i_{t}}$ with $i_{1}<\dots<i_{t},t\geq 2$, we define $N_{f}(w,a)=|\{p\in T(f)\mid\text{wt}(p_{[1,a]})=w\}|$ as the number of codewords with weight-$w$ in the first $a$ bits. Clearly, $P_{f}(w,a)=N_{f}(2^{m-t}-w,a)$. We have
@@
$N_{f}(w,a)=$  $\qquad\text{ if }a>2^{i_{t}}.$ (5d)
Here, $\alpha_{f}(s)=|\{1\leq j\leq t-1\mid i_{j}>s\}|$ is the number of elements in $\{i_{1},\dots,i_{t-1}\}$ that are larger than $s$, $f^{(0)}=x_{i_{1}}\cdots x_{i_{t-1}}$ and $f^{(s)}=x_{i_{1}}\cdots x_{i_{t-1}}x_{s}$, where $s\in B(f,t)$.
@@
[…]
@@
2) $2^{i_{t}-1}<a\leq 2^{i_{t}}$. From Lemma 4, there exists some automorphism $\varphi$, such that $p^{\prime}=\varphi(p)$ and it satisfies $p^{\prime}_{k}=p_{2^{i_{t}}+1-k}$ for $1\leq k\leq 2^{i_{t}}$. Therefore, $wt(p_{[1,a]})=wt(p^{\prime}_{[2^{i_{t}}+1-a,2^{i_{t}}]})=2^{i_{t}-t}-wt(p^{\prime}_{[1,2^{i_{t}}-a]})$. Since $\varphi$ is bijective, the number of codewords in $T(f)$ with $w$ ones in the first $a$ bits is equal to the number of codewords in $T(f)$ with $2^{i_{t}-t}-w$ ones in the first $2^{i_{t}}-a$ bits, which leads to
@@
Then $p_{[1,a]}=\rho(p)_{[1,a]}$ since $x_{i_{t}}=1$ for the first $2^{i_{t}-1}$ bits. Next, we will demonstrate that $\rho(p)\in T(g)$ for a suitable $g\preccurlyeq f$. The precise choice of $g$ depends on the value of $a_{i_{t},k}$ for $k\in B(f,t)$. This allows us to compute $N_{f}(w,a)$ iteratively. Consider the following three cases:
@@
III) If $a_{i_{t},k}=1$ for some $k\in B(f,t)$, then $\rho(p)\in T(f^{(s_{1})})$, where $s_{1}\in B(f,t)$ is the largest integer for which $a_{i_{t},s_{1}}=1$. In fact, several polynomials in $T(f)$ correspond to the same $p^{\prime}$ in $T(f^{(s_{1})})$, and this set is denoted as $H_{p^{\prime}}$. Specifically, $H_{p^{\prime}}\triangleq\{p\in T(f)\mid\rho(p)=p^{\prime}\}=\{p\in T(f)\mid p_{[1,2^{i_{t}-1}]}=(p^{\prime})_{[1,2^{i_{t}-1}]}\}$. Note that for any $p^{\prime}\in T(f^{(s_{1})})$ and $p^{\prime\prime}\in T(f^{(s_{2})})$, $(p^{\prime})_{[1,2^{i_{t}-1}]}\neq(p^{\prime\prime})_{[1,2^{i_{t}-1}]}$. Otherwise, by Lemma 3, $p^{\prime}=p^{\prime\prime}$ since $s_{1},s_{2}\leq i_{t}-1$, which is a contradiction. Therefore,
$H_{p^{\prime}}\cap H_{p^{\prime\prime}}=\varnothing.$ (6)
Next, we will prove that $|H_{p^{\prime}}|=2^{\alpha_{f}(s_{1})}$ for all $s_{1}\in B(f,t)$ and $p^{\prime}\in T(f^{(s_{1})})$.
Let $h=x_{s_{1}}+\sum_{k\in B(f,t),k<s}a_{i_{t},k}x_{k}+a_{i_{t},0}+1$ and $p^{\prime}=h\prod_{j=1}^{t-1}(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0})\in T(f^{(s_{1})})$. Now, for $i_{j}>s_{1}$, $h(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0})=h(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0}+h+1)$. Thus, for all $\beta_{j}\in\{0,1\}$, $h\prod_{i_{j}<s}(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0})\prod_{i_{j}>s,j\leq t-1}(x_{i_{j}}+\sum_{k\in B(f,j)}a_{i_{j},k}x_{k}+a_{i_{j},0}+\beta_{j}(h+1))$ are equal to $p^{\prime}$ in the first $2^{i_{t}-1}$ bits, so
$|H_{p^{\prime}}|\geq 2^{\alpha_{f}(s_{1})}.$ (7)
We have already shown there are at least $2^{\alpha_{f}(s_{1})}$ polynomials in $T(f)$ equal to $p^{\prime}$ in the first $2^{i_{t}-1}$ bits. Next we claim that $|H_{p^{\prime}}|$ is exactly $2^{\alpha_{f}(s_{1})}$. For $s_{1},s_{2}\in B(f,t)$, assume $s_{2}<s_{1}$ and there does not exist $s_{2}<s_{3}<s_{1}$ such that $s_{3}\in B(f,t)$. Then $\alpha_{f}(s_{1})+s_{1}=\alpha_{f}(s_{2})+s_{2}+1$. Furthermore, if $s_{0}$ is the minimum integer in $B(f,t)$, then $\alpha_{f}(s_{0})+s_{0}=t$. Hence, if $s$ is the $k$-thth smallest integer in $B(f,t)$, $\alpha_{f}(s)+s=t+k-1$. Since $|B(f,t)|=i_{t}-t$, we have
@@
$\lambda_{f}-2\lambda_{f^{(0)}}$ $\geq\sum_{s\in B(f,t)}\sum_{p^{\prime}\in T(f^{(s)})}|H_{p^{\prime}}|$
@@
Here, the first inequality is from (6). Note that the numbers of polynomials in case I) and II) are both $\lambda_{f^{(0)}}$, which needs to be subtracted. The second inequality is from (7). The equality is from multiplying both side of (8) by $\lambda_{f^{(0)}}$. Therefore, $|H_{p^{\prime}}|=2^{\alpha_{f}(s_{1})}$ for all $s_{1}\in B(f,t)$ and $p^{\prime}\in T(f^{(s_{1})})$.
@@
For $8\leq a\leq 16$, we have $P_{f}(w,a)=N_{f}(8-w,a)=N_{f}(6-w,a-8)=P_{f}(w+2,a-8)$, a result from (5d). Additionally, we note that $P_{f}(w,a)=P_{f}(14-w,8-a)$ when $4\leq a\leq 8$ according to (5c).
For $1\leq a\leq 4$ and $w\neq 8$, $P_{f}(w,a)$ is calculated by (5). In fact, we know that a minimum-weight codeword in $T(x_{2}x_{3})$ can be represented as
@@
When $1\leq a\leq 4$, $x_{3}=1$, so $p_{[1,4]}=p^{\prime}_{[1,4]}$, where
$p^{\prime}=(a_{3,1}x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}x_{1}+a_{2,0}).$
We consider three cases for $p^{\prime}$:
Case I: $a_{3,1}=0,a_{3,0}=1$, then $p^{\prime}=0$.
Case II: $a_{3,1}=a_{3,0}=0$, then $p^{\prime}=x_{2}+a_{2,1}x_{1}+a_{2,0}\in T(x_{2})$.
Case III: $a_{3,1}=1$, then $p^{\prime}=(x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}x_{1}+a_{2,0})\in T(x_{2}x_{1})$. Note that $(x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}x_{1}+a_{2,0})=(x_{1}+a_{3,0}+1)(x_{2}+a_{2,1}(a_{3,0}+1)+a_{2,0})$. One polynomial $p^{\prime}$ in $T(x_{2}x_{1})$ corresponds to two polynomials in $T(x_{1})$.
In conclusion, we have $N_{f}(w,a)=N_{x_{2}}(w,a)+2N_{x_{1}x_{2}}(w,a)$ for $1\leq a\leq 4$, and $P_{f}(w,a)=N_{f}(8-w,a)$. Finally, $P_{f}(8,a)=N_{f}(0,a)=16-\sum_{w=1}^{8}N_{f}(w,a)$ according to (5).
@@ IV-B Number of minimum-weight codewords in QUP polar codes
Building upon Lemma 5, we derive the weight spectrum of $T_{P}(f,X_{i})$ after puncturing. Next, we prove that a minimum-weight codeword after puncturing must also be a minimum-weight codeword at the mother code length, that is, it belongs to $T_{P}(f,X_{i})$.
[…]
@@
$({\bm{f}}_{N}^{(j)})_{[N]/X_{i}}$ is a minimum-weight codeword of $\mathcal{C}(m,(\bm{0}_{1}^{j-1},1),X_{i})$. Moreover, the minimum-weight codewords of $\mathcal{C}(m,(\bm{0}_{1}^{j-1},1),X_{i})$ are also minimum-weight codewords at the mother code length.
@@
The proof is by induction on $m$. The base case, $m=1$, can be proved directly. For the inductive step, from m1m-1 to $m$, we consider two cases:
1) $1\leq j\leq N/2$, which means ${\bm{f}}_{N}^{(j)}$ is in the top half of $\bm{F}_{N}$. Then $\mathcal{C}(m,(\bm{0}_{1}^{j-1},1),X_{i})=\{(\bm{u}_{[N/2]/X_{i}}\oplus\bm{v}_{[N/2]/X_{i}},\bm{v})\mid\bm{u}\in\mathcal{C}(m-1,(\bm{0}_{1}^{j-1},1)),\bm{v}\in\mathbb{F}_{2}^{N/2}\}$. Then by the induction hypothesis,
@@
Furthermore, if $(\bm{u}_{[N/2]/X_{i}}+\bm{v}_{[N/2]/X_{i}},\bm{v})$ is a minimum-weight codeword, then the equality in (IV-B) must hold. By inductive hypothesis, $\text{wt}(\bm{u}_{[N/2]/X_{i}})=\text{wt}(({\bm{f}}_{N/2}^{(j)})_{[N/2]/X_{i}})$ implies that $\bm{u}$ is the minimum-weight codeword at the mother code length $N/2$. In addition, $\text{wt}((\bm{u}_{[N/2]/X_{i}}\oplus\bm{v}_{[N/2]/X_{i}},\bm{v}))=\text{wt}(\bm{u}_{[N/2]/X_{i}})$ implies that $\text{wt}(\bm{v}_{X_{i}})=0$ and $\bm{v}_{[N/2]/X_{i}}$ is covered by $\bm{u}_{[N/2]/X_{i}}$, i.e., $v_{k}=1$ implies $u_{k}=1$ for any $k\in[N/2]/X_{i}$. Hence, $\bm{v}$ is covered by $\bm{u}$. Then $\text{wt}((\bm{u}\oplus\bm{v},\bm{v}))=\text{wt}(\bm{u})=\text{wt}({\bm{f}}_{N/2}^{(j)})=\text{wt}({\bm{f}}_{N}^{(j)})$, so $(\bm{u}\oplus\bm{v},\bm{v})$ is the minimum-weight codeword at the mother code length $N$.
@@
Furthermore, if $(\bm{v}_{[N/2]/X_{i}},\bm{v})$ is the minimum-weight codeword, then the equality in (IV-B) must hold, that is, $\text{wt}(\bm{v})=\text{wt}({\bm{f}}_{N/2}^{(j)})$. Then $(\bm{v},\bm{v})$ is the minimum-weight codeword at the mother code length $N$.
In conclusion, the minimum-weight codewords in $C(\mathcal{I},X_{i})$ are also minimum-weight codewords at the mother code length, meaning they belong to some $T_{P}(f,X_{i})$. ∎
The number of minimum-weight codewords in QUP polar codes follows directly from Lemma 5 and Lemma 6.
@@
Let $C(\mathcal{I},X_{i})$ be a punctured decreasing $r$-th order polar code with length $N=2^{m}$ and puncturing pattern $X_{i}$, where $X_{i}$ is the first $i$ bits of the sequence $\bm{q}$. Denote $d$ to be the minimum positive integer such that $\sum_{f\in\mathcal{I}}P_{f}(d,X_{i})>0$. Then $d$ is the minimum weight of $C(\mathcal{I},X_{i})$ and $A_{d}(C_{P}(\mathcal{I},i))=\sum_{f\in\mathcal{I}}P_{f}(2^{m-r}-d,X_{i})$ and $A_{w}(C_{P}(\mathcal{I},i))\geq\sum_{f\in\mathcal{I}}P_{f}(2^{m-r}-w,X_{i})$ for $w>d$.
The Algorithm 2 concludes the procedure for calculating the number of minimum codewords in QUP polar code. It is worth noting that the lower bound $\sum_{f\in\mathcal{I}}P_{f}(2^{m-r}-w,i)$ in Theorem 3 for $w$ greater than the minimum weight can also be calculated using Lemma 5.
Algorithm 2 Calculate the number of minimum-weight codewords in QUP polar code
@@
As the code length $N$ approaches infinity, the time complexity of Algorithm 2 is $O(N^{3})$.
@@
Since an LTA transformation is an automorphism of decreasing polar codes, applying this transformation to the puncturing pattern will result in the same weight spectrum. Consequently, we can determine the weight spectrum for a series of punctured decreasing polar codes.